# Active Sensing of Knee Osteoarthritis Progression with Reinforcement Learning

### v2

@@ 2.3 Proposed method
For each subject, at time point $t$, an agent takes one of two actions – to follow up or to dismiss (i.e. skip examination) at time $t+1$. The rewards for these actions are denoted as $r_{f}$ and $r_{d}$, respectively. The reference point $t_{r}$ is updated to the latest visit, ensuring that the reward calculations are based on the most recent state of the subject’s condition. By comparing $t+1$ with $t_{p}$ – the nearest progression time point $t_{p}$ extracted from the patient-level progressions set $\mathbf{M}_{p}$, we derived five possible scenarios. Namely, there are three cases for the correctness of the follow-up action ($a=1$): early visit, timely visit, and late visit and two cases for the one of the dismissal action ($a=0$): true dismissal and false dismissal.
@@ 3.1 Knee osteoarthritis
The main limitation of most previous studies on KOA progression prediction is that they consider predictions per knee, disregarding that OA often co-occurs in two knees. The second limitation is that they ignore the long disease progression timelines, taking for modeling only one baseline measurement point. OA can progress very slowly, and be drastically accelerated by various future unknown factors, e.g. injury [poulsen2019knee]. Therefore, we believe that it is crucial to update the model’s predictions when newer data from a patient becomes available. Finally, none of the previous studies considered the financial aspect of their results, nor downstream drug development or clinical applications. Our work aims to address all these limitations, enabling dynamic patient-level data collection, while balancing data collection utility and costs with RL.
@@ 3.2 Active sensing
RL has become one of the most common techniques to explore AS in recent years, particularly in medical applications for disease management [2, 83, 10, 62, 29]. DPSCREEN framework proposed by [2] aimed to deliver a personalized screening policy for breast cancer (BC) patients. The proposed policy takes into account both patient features and their clinical history, and it demonstrated significant reductions in the number of screenings performed while maintaining the same expected delays in disease detection. A similar problem of BC screening was also later tackled by [81].
@@ 5.2 Impact of financial factors on decision-making
To gain deeper insights into the trade-offs related to the cost parameters $\lambda$ and $c$, and their impact on the outcomes, we investigated the relationship between the relative cost $\frac{\lambda}{c}$ and the resulting RPP normalized over the hospital cost, $\frac{\bar{R}}{c}$. Specifically, we varied the hospital visit cost parameter, $\lambda$, across the range $\{0,0.25,0.5,0.75,1,1.5,2,2.5,3,3.5,4\}$, while adjusting the MLAC, $c$, within the set $\{0.5,1,1.5,2,3,4\}$. The association is illustrated in Figure 4c. We observed that $\frac{\bar{R}}{c}$ exhibited an inverse, non-linear relationship with the increasing of $\frac{\lambda}{c}$. In other words, decreasing $\frac{\lambda}{c}$ was beneficial for the resulting RPP. Specifically, when the ratio $\frac{\lambda}{c}$ declines to $0.3$, it becomes the critical trade-off point at which our policy begins to yield positive RPPs.
@@ 5.3 Patient-level versus knee-level predictions
Table 4 demonstrates the overall performance of CLIMATv2 and our proposed method with knee-level and patient-level approaches. Notably, the patient-level approach outperformed the knee-level approach for both CLIMATv2 and the RL-based policy, with average improvements in BA of $1.44\%$ and $3.73\%$, respectively, and in RC of $4.58\%$ and $20.25\%$, respectively. Both the methods gained a $0.02$ increase in RPP when switching from the knee-level to the patient-level approach.
@@ 5.7 Follow-up frequency
In this section, we assessed the dependency between the number of follow-up decisions made by our method and the disease characteristics. Specifically, Figure 8 visualizes the number of hospital visits per person according to individual KLG pairs measured at the baseline visit. Those KLG pairs are classified into $4$ subgroups, denoted by color regions: ”Healthy” when neither knee has KLG greater than 1, ”Mild” when neither knee has KLG greater than 2, ”Moderate” when neither knee has KLG greater than 3, and ”Severe” when at least one knee has KLG 4. We illustrated our policies with $3$ different $\lambda$ of $0.5$, $0.8$, and $1.0$ by three lines in the graphs. All these policies delivered at least one follow-up to participants regardless of the costs. The diagram shows that our policy learned to be severity-aware, and realized that patients with more severe forms of OA in both knees progress faster. That is, the agent recommended more visits in such a case, and consistently in all considered cost settings.
@@ 6 Discussion and conclusions
KOA is so far a non-reversible joint disease. The common trait of KOA progression is the narrowing of JSW, which is observable via various modalities. To detect and track KOA progression longitudinally, imaging techniques like X-rays, MRI, and CT scans are often employed [67]. We proposed to connect quantitatively measured radiographic changes with data acquisition cost, aiming to define the utility of detecting an KOA progression event. Although quality-of-life (QoL) measures can be more informative disease markers to track, the cause of pain at different stages of KOA is different [66], the relationship between pain, joint function, and imaging can change, thereby affecting the predictive model. Moreover, it is challenging to measure QoL objectively due to chronic pain having a psychosocially driven nature  [71]. Hence, in our study, we concluded that the best available method to track the disease in OA is based on radiographs due to their wide accessibility, and ease of use.
@@
In this work, we developed a reward function, which balances the ability of agent to detect KOA progression and the acquisition costs. The proposed reward incorporates follow-up and dismissal actions, divided into mutually exclusive scenarios that occur during following up the disease progression in a patient. However, we faced challenges in defining the value of true dismissal and late follow-up decisions as the disease statuses from those cases are unknown. To address the challenges, we introduced hyper-parameters $\alpha$ and $\beta$ into our reward function. $\alpha$ mitigates the punishment for late follow-up over false dismissal (late better than none) and $\beta$ represents the gained reward of true dismissal. We believe that $\alpha$ and $\beta$ should be tailored to better fit different socio-economic contexts and applications, investigating which is beyond the scope of the present work.

### v3

@@ 2.3 Proposed method
For each subject, at time point $t\geq 0$, an agent takes one of two actions – to follow up or to dismiss (i.e. skip examination) at time $t+1$. The rewards for these actions are denoted as $r_{f}$ and $r_{d}$, respectively. The reference point $t_{r}$ is updated to the latest visit, ensuring that the reward calculations are based on the most recent state of the subject’s condition. By comparing $t+1$ with $t_{p}$ – the nearest progression time point $t_{p}$ extracted from the patient-level progressions set $\mathbf{M}_{p}$, we derived five possible scenarios. Namely, there are three cases for the correctness of the follow-up action ($a=1$): early visit, timely visit, and late visit and two cases for the one of the dismissal action ($a=0$): true dismissal and false dismissal.
@@ 3.1 Knee osteoarthritis
The main limitation of most previous studies on KOA progression prediction is that they consider predictions per knee, disregarding that OA often co-occurs in two knees. The second limitation is that they ignore the long disease progression timelines, taking only the baseline measurement point for modeling. OA can progress very slowly, and be drastically accelerated by various future unknown factors, e.g. injury [poulsen2019knee]. Therefore, we believe that it is crucial to update the model’s predictions when newer data from a patient becomes available. Finally, none of the previous studies considered the financial aspect of their results, nor downstream drug development or clinical applications. Our work aims to address all these limitations, enabling dynamic patient-level data collection, while balancing data collection utility and costs with RL.
@@ 3.2 Active sensing
RL has become one of the most common techniques to explore AS in recent years, particularly in medical applications for disease management [2, 83, 10, 62, 29]. DPSCREEN framework proposed by [2] aimed to deliver a personalized screening policy for breast cancer (BC) patients. The proposed policy utilizes both patient features and their clinical history, and it demonstrated significant reductions in the number of screenings performed while maintaining the same expected delays in disease detection. A similar problem of BC screening was also later tackled by [81].
@@ 5.2 Impact of financial factors on decision-making
To gain deeper insights into the trade-offs related to the cost parameters $\lambda$ and $c$, and their impact on the outcomes, we investigated the relationship between the relative cost $\frac{\lambda}{c}$ and the resulting RPP normalized over the hospital cost, $\frac{\bar{R}}{c}$. Specifically, we varied the hospital visit cost parameter, $\lambda$, across the range $\{0,0.25,0.5,0.75,1,1.5,2,2.5,3,3.5,4\}$, while adjusting the MLAC, $c$, within the set $\{0.5,1,1.5,2,3,4\}$. The association is illustrated in Figure 4c. We observed that $\frac{\bar{R}}{c}$ exhibited an inverse, non-linear relationship with the increasing of $\frac{\lambda}{c}$. In other words, decreasing $\frac{\lambda}{c}$ was beneficial for the resulting RPP. Specifically, when the ratio $\frac{\lambda}{c}$ declines to $0.3$, it becomes the critical trade-off point at which our policy begins to yield positive RPPs. The raw experimental data behind Figure 4c is shown in Suppl. Table S2
@@ 5.3 Patient-level versus knee-level predictions
Table 4 demonstrates the overall performance of CLIMATv2 and our proposed method with knee-level and patient-level approaches (see Suppl. Table S1 for raw data). Notably, the patient-level approach outperformed the knee-level approach for both CLIMATv2 and the RL-based policy, with average improvements in BA of $1.44\%$ and $3.73\%$, respectively, and in RC of $4.58\%$ and $20.25\%$, respectively. Both the methods gained a $0.02$ increase in RPP when switching from the knee-level to the patient-level approach.
@@ 5.7 Follow-up frequency
In this section, we assessed the dependency between the number of follow-up decisions made by our method and the disease characteristics. Specifically, Figure 8 visualizes the number of hospital visits per person according to individual KLG pairs measured at the baseline visit (see Suppl. Table S3 for raw data). Those KLG pairs are classified into $4$ subgroups, denoted by color regions: ”Healthy” when neither knee has KLG greater than 1, ”Mild” when neither knee has KLG greater than 2, ”Moderate” when neither knee has KLG greater than 3, and ”Severe” when at least one knee has KLG 4. We illustrated our policies with $3$ different $\lambda$ of $0.5$, $0.8$, and $1.0$ by three lines in the graphs. All these policies delivered at least one follow-up to participants regardless of the costs. The diagram shows that our policy learned to be severity-aware, and realized that patients with more severe forms of OA in both knees progress faster. That is, the agent recommended more visits in such a case, and consistently in all considered cost settings.
@@ 6 Discussion and conclusions
KOA is so far a non-reversible joint disease. The common trait of KOA progression is the narrowing of JSW, which is observable via various modalities. To detect and track KOA progression longitudinally, imaging techniques like X-rays, MRI, and CT scans are often employed [67]. We proposed to connect quantitatively measured radiographic changes with data acquisition cost, aiming to define the utility of detecting a KOA progression event. Although quality-of-life (QoL) measures can be more informative disease markers to track, the cause of pain at different stages of KOA is different [66], the relationship between pain, joint function, and imaging can change, thereby affecting the predictive model. Moreover, it is challenging to measure QoL objectively due to chronic pain having a psychosocially driven nature  [71]. Hence, in our study, we concluded that the best available method to track the disease in OA is based on radiographs due to their wide accessibility, and ease of use.
@@
In this work, we developed a reward function, which balances the ability of the agent to detect KOA progression and the acquisition costs. The proposed reward incorporates follow-up and dismissal actions, divided into mutually exclusive scenarios that occur during the follow-up of the disease progression in a patient. However, we faced challenges in defining the value of true dismissal and late follow-up decisions as the disease statuses from those cases are unknown. To address the challenges, we introduced hyper-parameters $\alpha$ and $\beta$ into our reward function. $\alpha$ mitigates the punishment for late follow-up over false dismissal (late better than none) and $\beta$ represents the gained reward of true dismissal. We believe that $\alpha$ and $\beta$ should be tailored to better fit different socio-economic contexts and applications, investigating which is beyond the scope of the present work.
@@
S1.2 Data behind figures
Table S1: Balanced accuracy and recall scores (mean and standard error, %) of our method and CLIMAT performed with knee-level and patient-level settings, presented at each year. Bold numbers represent the highest scores among all methods.
Method Year(s) BA(%) RC(%)
Knee-level CLIMATv2 1 57.09 $\pm$ 1.69 24.29 $\pm$ 6.22
2 50.87 $\pm$ 1.26 23.77 $\pm$ 4.30
3 56.34 $\pm$ 0.84 46.15 $\pm$ 5.00
4 55.60 $\pm$ 1.16 38.59 $\pm$ 4.55
Ours 1 60.43 $\pm$ 0.72 58.98 $\pm$ 3.84
2 59.26 $\pm$ 0.57 82.92 $\pm$ 1.83
3 47.30 $\pm$ 0.80 32.88 $\pm$ 0.72
4 59.41 $\pm$ 0.73 38.51 $\pm$ 1.95
Patient-level CLIMATv2 1 51.83 $\pm$ 0.77 7.55 $\pm$ 3.18
2 56.71 $\pm$ 0.95 42.83 $\pm$ 5.97
3 58.73 $\pm$ 0.79 50.58 $\pm$ 3.24
4 58.38 $\pm$ 0.89 50.16 $\pm$ 4.80
Ours 1 62.81 $\pm$ 0.57 74.08 $\pm$ 2.62
2 53.56 $\pm$ 0.44 96.98 $\pm$ 0.85
3 60.40 $\pm$ 0.67 67.69 $\pm$ 1.03
4 64.55 $\pm$ 0.74 55.94 $\pm$ 1.56
Table S2: RPP achieved by baselines and our policy from various cost settings, where (a) adjusting follow-up cost $\lambda$ while keeping constant MLAC $c=2$, (b) hold $\lambda=0.5$ while varying $c$, (c) performance of our policy on grid combination between $\lambda$ and $c$.
$\lambda$ NS ANS CLIMATv2 Ours
0.00 -1.64 1.14 -0.51 1.18
0.25 -1.64 -0.06 -0.74 0.54
0.50 -1.64 -0.86 -0.89 0.34
0.75 -1.64 -2.06 -1.12 -0.23
1.00 -1.64 -2.86 -1.27 -0.69
1.50 -1.64 -4.86 -1.65 -1.10
2.00 -1.64 -6.86 -2.03 -1.16
(a) With $c=2$, varying $\lambda$
$c$ NS ANS CLIMATv2 Ours
0.50 -0.41 -1.71 -0.51 -0.36
1.00 -0.82 -1.43 -0.63 -0.35
1.50 -1.23 -1.15 -0.76 -0.05
2.00 -1.64 -0.87 -0.89 0.34
(b) With $\lambda=0.5$, varying $c$
$\lambda$ $c$
0.50 1.00 1.50 2.00 3.00 4.00
0.00 0.31 0.61 0.92 1.18 1.85 2.45
0.25 -0.14 0.15 0.36 0.54 1.19 1.83
0.50 -0.36 -0.36 -0.05 0.34 0.65 1.20
0.75 -0.44 -0.54 -0.48 -0.23 0.16 0.71
1.00 -0.49 -0.65 -0.75 -0.69 -0.09 0.16
1.50 -0.45 -0.83 -0.97 -1.10 -1.01 -0.52
2.00 -0.42 -0.90 -1.16 -1.16 -1.49 -1.35
2.50 -0.41 -0.85 -1.31 -1.45 -1.61 -1.85
3.00 -0.41 -0.85 -1.22 -1.75 -1.68 -2.09
3.50 -0.41 -0.83 -1.21 -1.78 -2.01 -2.18
4.00 -0.41 -0.82 -1.28 -1.78 -2.22 -2.33
(c) RPP gained by our RL policy depending on $c$ and $\lambda$ combinations. Underscored numbers indicate positive or close to 00 RPP values.
Table S3: Number of visits (mean and standard error) of $3$ policies according to pairs of KLGs at baseline. Those policies were trained with different follow-up costs hyper-parameter $\lambda=\{0.5,0.8,1.0\}$.
Baseline KLGs Polices
$\lambda=0.5$ $\lambda=0.8$ $\lambda=1.0$
(0, 0) 1.33 $\pm$ 0.18 1.07 $\pm$ 0.10 1.03 $\pm$ 0.06
(0, 1) 1.34 $\pm$ 0.20 1.13 $\pm$ 0.11 1.04 $\pm$ 0.07
(1, 1) 1.65 $\pm$ 0.29 1.31 $\pm$ 0.22 1.14 $\pm$ 0.13
(0, 2) 1.94 $\pm$ 0.32 1.46 $\pm$ 0.26 1.18 $\pm$ 0.15
(1, 2) 2.01 $\pm$ 0.33 1.52 $\pm$ 0.23 1.31 $\pm$ 0.18
(2, 2) 2.06 $\pm$ 0.33 1.59 $\pm$ 0.24 1.25 $\pm$ 0.16
(0, 3) 3.03 $\pm$ 0.31 2.31 $\pm$ 0.30 1.51 $\pm$ 0.20
(1, 3) 2.83 $\pm$ 0.32 2.20 $\pm$ 0.30 1.53 $\pm$ 0.20
(2, 3) 2.96 $\pm$ 0.32 2.16 $\pm$ 0.29 1.76 $\pm$ 0.26
(3, 3) 3.61 $\pm$ 0.21 2.88 $\pm$ 0.28 2.15 $\pm$ 0.25
(1, 4) 3.83 $\pm$ 0.12 3.30 $\pm$ 0.23 2.02 $\pm$ 0.21
(2, 4) 3.38 $\pm$ 0.17 2.88 $\pm$ 0.25 2.28 $\pm$ 0.23
(3, 4) 3.96 $\pm$ 0.06 3.24 $\pm$ 0.28 2.24 $\pm$ 0.27